# Graph-SCP: Accelerating Set Cover Problems with Graph Neural Networks

### v2

@@ Appendix A Appendix
GNN Architecture Study
We compare 4 commonly used GNN architectures: Graph Attention Networks (GAT) (Veličković et al. 2018), Graph Convolutional Network (GCN) (Kipf and Welling 2017), Graph Isomorphism Networks (GIN) (Xu et al. 2018) and Chebyshev GCN (Defferrard, Bresson, and Vandergheynst 2016). We designed each of these models to have similar number of parameters and were trained on the same train set and tested on Instance Type 1, with results shown in Table 4. We observe that GraphSAGE and Chebyshev GCN outperform other architectures, with GraphSAGE achieving the highest speedup at approximately $8$x. A potential source of the differences in performance could be attributed to each model being limited to a similar number of training parameters (approximately $2$ million).
GNN Type Speedup Factor
GraphSAGE $8.20\pm 1.17$
Chebyshev $7.98\pm 0.74$
GIN $4.16\pm 0.68$
GCN $3.74\pm 0.92$
GAT $3.52\pm 0.51$
Table 4: Speedup factors across various GNN architectures. We constructed all GNN models to have similar number of parameters (approximately $2$ million). We observe that GraphSAGE and Chebyshev GCN outperform other architectures, with GraphSAGE achieving the highest speedup at approximately $8$x. Results were generated from Instance Type 1, but are representative across instances.
Hardware
We train and test on an Intel Xeon Gold 6148 CPU with 24 cores and one NVIDIA Tesla V100 GPU
Reproducibility
Our code repository is available at https://github.com/zohairshafi/Graph-SCP

### v3

@@ 5.1 Discussions & Limitations
Table 4: Speedup factors when training Graph-SCP without the unsupervised component result in lower acceleration. Further removal of hypergraph node features leads to an additional decline in acceleration.
Method Type 1 Type 2 Type 3 Type 4
Graph-SCP $9.97\pm 2.13$ $70.148\pm 18.27$ $11.76\pm 2.43$ $3.20\pm 0.84$
w/o Unsupervised $6.75\pm 3.43$ $64.59\pm 22.00$ $5.83\pm 0.56$ $2.23\pm 0.32$
w/o Unsupervised
& Hypergraph
$4.45\pm 1.73$ $60.869\pm 18.85$ $3.54\pm 0.45$ $2.014\pm 0.427$
Finally, we also compare Graph-SCP’s performance by replacing Graph-SAGE as the GNN used with Graph Attention Networks (GAT) [36], Graph Convolutional Network (GCN) [22], Graph Isomorphism Networks (GIN) [39], and Chebyshev GCN [12] with average runtime speed-ups of $3.52\pm 0.51$, $3.74\pm 0.92$, $4.16\pm 0.68$ and $7.98\pm 0.74$, respectively, on Instance Type 1. We posit that differences in performance are due to limiting the sizes of each model to approximately 2 million parameters to be comparable to each other.
6 Related Work and Conclusions
Existing literature explores the use of ML to replace or augment traditional solvers for various CO problems with the objective of accelerating runtimes and, where appropriate, improving the quality of solutions. The proposed methods can be classified into two approaches: learning a model in an end-to-end manner to replace a CO solver, or learning a model to aid an existing CO solver.
Under the category that uses ML to replace a solver, Numeroso et al. [28] apply the Neural Algorithmic Reasoning framework proposed by Veličković et al. [35] to jointly optimize for the primal and dual of a given problem. They train models to find joint solutions for both the min-cut and max-flow problems and show that their solution leads to better overall performance. For NP-hard problems, Khalil et al. [11] propose a reinforcement learning approach that uses graph embeddings to learn a greedy policy that incrementally builds a solution. Heydaribeni et al. [19] model constrained CO problems as a hypergraph followed by mapping continuous values generated by the Hypergraph Neural Network to integer node assignments using simulated annealing. Schuetz et al. [30] and Hu et al. [20] use GNNs to minimize the Quadratic Unconstrained Binary Optimization (QUBO) formulations of CO problems. Boisvert et al. [5] provide a generic method for encoding arbitrary CO problems into a graph structure for use with GNNs. Yau et al. [40] propose GNNs that can capture provably optimal message passing algorithms for a large class of combinatorial optimization problems. The aim of our work is to focus on accelerating an existing solver, allowing us to carry over its performance guarantees. In a similar context, Ireland et al. [21] and Tian et al. [33] look at problems like minimum vertex cover and max-cut and use a GNN to identify promising nodes for large scale graphs. Han et al. [18] train a GNN to predict a marginal probability of each variable and then construct a trust region to search for high quality feasible solutions. Verhaeghe et al. [37] cast the resource constrained project scheduling problem as a directed graph and use GNNs to extract information to help speed up a traditional solver. Li et al. [25] use a GNN to predict if a vertex is part of the optimal solution followed by a tree search to arrive at the final solution. However, Bother et al. [6] later showed that the results in [25] are not reproducible. Liu et al. [26] investigate the performance of branching rules with respect to the size of the search neighborhood and devise a framework for guiding the search of the branching heuristic. Ding et al. [13] also use a learned GCN model to predict branching cuts. They modify the solver by adding a local branching rule in their approximate case and performing an actual branch at the root node in the exact case. Nair et al. [27] propose two methods, Neural Branching and Neural Diving. While we operate exclusively on the input space irrespective of a solver, neural diving aims to explore the branch and bound tree in a depth-first manner. They differ from traditional diving in that they do not start from the root and explore all the way to the leaf node, but start from root, stop mid way and allow a solver to solve the remaining subproblem. Our method similarly uses a traditional solver to find a solution. Gasse et al. [14] use a GNN to learn computationally expensive branch-and-bound heuristics for MIP solvers. The learning is done offline and the learned models are then used at runtime in place of the heuristic to achieve faster runtimes while maintaining the quality of the solution. This work is similar to our approach in that it uses ML to aid a CO solver; however, it does so by speeding up the internal workings (branching heuristic) of the solver. These methods lie on the lines of modifying or improving the branch and bound heuristic. Kruber et al. [23] use supervised learning to determine when a Dantzig-Wolfe decomposition should be applied to a Mixed Integer Program (MIP). Graph-SCP differs in that it reduces the input problem size complexity, and therefore can be used in conjunction with any existing solver, while achieving the optimal solution. Shen et al. [32] use support vector machines to speed up column generation methods by predicting the solution to the pricing problem at each step. We point the interested reader to surveys by Bengio et al. [3] and Cappart et al. [7] that provide a detailed summary of research in this field.
Our contribution is Graph-SCP, a GNN framework to accelerate runtimes for both conventional and ML augmented CO solvers. By identifying a smaller subproblem that encapsulates the solution space, Graph-SCP achieves runtimes up to approximately 1010x faster, while maintaining the reliability associated with traditional solvers. A distinguishing feature of Graph-SCP is its GNN prediction module, which is executed only once. This efficiency is in contrast to many ML solutions for CO problems. We illustrate Graph-SCP’s robust generalization across SCP instances with diverse characteristics and complexities, showing its effectiveness on instances from the canonical OR Library. Through comparative studies, we offer insights into the impact of different modeling choices on Graph-SCP’s performance. Crucially, since Graph-SCP operates only in the input space, it seamlessly integrates with any traditional or ML-augmented CO solver.
{credits}
6.0.1 Acknowledgments.
Distribution Statement A. Approved for public release. Distribution is unlimited. This material is based upon work supported by the Under Secretary of Defense for Research and Engineering under Air Force Contract No. FA8702-15-D-0001. Any opinions, findings, conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the Under Secretary of Defense for Research and Engineering. © 2025 Massachusetts Institute of Technology. Delivered to the U.S. Government with Unlimited Rights, as defined in DFARS Part 252.227-7013 or 7014 (Feb 2014). Notwithstanding any copyright notice, U.S. Government rights in this work are defined by DFARS 252.227-7013 or DFARS 252.227-7014. Use of this work other than as specifically authorized by the U.S. Government may violate any copyrights that exist in this work.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.